the speeding car is going 68 miles per hour. the exit is in 3.4 miles. officer tries to start her motorcycle it wont start. finally thirty seconds later it starts and she is in persuite of the speeding car. write an expression showing the time the chasing officer has been traveling. and what speed the officer will have to average to catch up to the speeding car before the exit
Answer by lwsshak3(11628)

the speeding car is going 68 miles per hour. the exit is in 3.4 miles. officer tries to start her motorcycle it wont start. finally thirty seconds later it starts and she is in persuite of the speeding car. write an expression showing the time the chasing officer has been traveling. and what speed the officer will have to average to catch up to the speeding car before the exit
***
let x=catch-up speed of officer
travel time of speeding car=distance/speed=(3.4/68) hr
travel time of officer=(3.4/68) hr-30 sec=(3.4/68)-(1/120)≈.042 hr
.042x=3.4
x≈81.6 mph
catch-up speed of officer≈81.5 mph